OpenStudy (anonymous):

p=2l+2w slove for L P is 50 With is (8x-15) Help me slove. Thanks:)

OpenStudy (anonymous):

Linear Programming (LP)?

OpenStudy (anonymous):

? maybe a typo here?

OpenStudy (agreene):

the way I understand what you are saying is this: p=2L+2w with p=50 And W=(8x-15) If so, we can plug in for p and w and get 50=2L+2(8x-15) solving for L: start by distributing the 2 50=2L+16x-120 add 120 and subtract 16x from both sides 170-6x=2L divide by 2 (170-6x)/2 = L 85-3x=L